OpenStudy (aaronandyson):

(1+tan^2(A) times cot(A )all over cosec^2(A) = tan(A)

OpenStudy (aaronandyson):

@Michele_Laino

OpenStudy (michele_laino):

here, a possible way, is to replace tanA with sinA/cosA, and cosecA with 1/sinA, in so doing the left side becomes: \[\Large \begin{gathered} \frac{{{{\left( {1 + \tan A} \right)}^2}\cot A}}{{{{\left( {\csc A} \right)}^2}}} = \hfill \\ \hfill \\ = \frac{{{{\left( {1 + \frac{{\sin A}}{{\cos A}}} \right)}^2}\frac{{\cos A}}{{\sin A}}}}{{{{\left( {\frac{1}{{\sin A}}} \right)}^2}}} = ... \hfill \\ \end{gathered} \]

OpenStudy (aaronandyson):

The question is \[(1 + \tan^2(A).\cot(A) \over cosec^2(A) \] is equal to tan(A)

OpenStudy (michele_laino):

yes! it is an identity, so we have to show that the left side is equal to the right side

OpenStudy (aaronandyson):

so sec^2(A).cot(A) over cosec^2(A)

OpenStudy (michele_laino):

sorry I have made an error, here are your right steps: \[\Large \begin{gathered} \frac{{\left( {1 + {{\left( {\tan A} \right)}^2}} \right)\cot A}}{{{{\left( {\csc A} \right)}^2}}} = \hfill \\ \hfill \\ = \frac{{\left( {1 + {{\left( {\frac{{\sin A}}{{\cos A}}} \right)}^2}} \right)\frac{{\cos A}}{{\sin A}}}}{{{{\left( {\frac{1}{{\sin A}}} \right)}^2}}} = ... \hfill \\ \end{gathered} \]

OpenStudy (aaronandyson):

sec^2(A)

OpenStudy (michele_laino):

hint: here are more steps: \[\Large \begin{gathered} \frac{{\left( {1 + {{\left( {\tan A} \right)}^2}} \right)\cot A}}{{{{\left( {\csc A} \right)}^2}}} = \hfill \\ \hfill \\ = \frac{{\left( {1 + {{\left( {\frac{{\sin A}}{{\cos A}}} \right)}^2}} \right)\frac{{\cos A}}{{\sin A}}}}{{{{\left( {\frac{1}{{\sin A}}} \right)}^2}}} = \hfill \\ \hfill \\ = \frac{{{{\left( {\sin A} \right)}^2} + {{\left( {\cos A} \right)}^2}}}{{{{\left( {\cos A} \right)}^2}}} \cdot \frac{{\cos A}}{{\sin A}} \cdot {\left( {\sin A} \right)^2} = ... \hfill \\ \end{gathered} \]

OpenStudy (aaronandyson):

didn't get the last step

OpenStudy (michele_laino):

we have the last step because we can write this: \[\Large \left( {1 + {{\left( {\frac{{\sin A}}{{\cos A}}} \right)}^2}} \right) = \frac{{{{\left( {\sin A} \right)}^2} + {{\left( {\cos A} \right)}^2}}}{{{{\left( {\cos A} \right)}^2}}}\]

OpenStudy (michele_laino):

furthermore, we can write this: \[\Large \frac{1}{{{{\left( {\frac{1}{{\sin A}}} \right)}^2}}} = \frac{1}{{\frac{1}{{{{\left( {\sin A} \right)}^2}}}}} = {\left( {\sin A} \right)^2}\]

OpenStudy (aaronandyson):

1/cos^2(A) . cos(A)/sin(A) . (sin^2(A))

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

now continue to simplify

OpenStudy (aaronandyson):

confused ;-;

OpenStudy (michele_laino):

\[\Large \frac{1}{{{{\left( {\cos A} \right)}^2}}} \cdot \frac{{\cos A}}{{\sin A}} \cdot {\left( {\sin A} \right)^2} = \frac{{\sin A}}{{\cos A}}\]